OpenStudy (anonymous):

U^2 - 2u +? equals a perfect square

OpenStudy (anonymous):

Take 1/2 of the u term and square that number. That will replace the ? and you will have a perfect square.

OpenStudy (anonymous):

so 2 would be the ?

OpenStudy (anonymous):

No. I said take 1/2 of the u term, then square that number. The coefficient of the u term is -2. So take 1/2 of -2, which is -1. Now square -1, you get (-1)^2 = +1 So adding +1 will make it a perfect square, as you will have u^2 - 2u + 1 = (u-1)(u-1) = (u-1)^2.